OpenStudy (anonymous):

2g+14+5g=8g-3+8 yes i know another question haha im officially dumb but please help

OpenStudy (anonymous):

put all like terms on each side of equation

OpenStudy (anonymous):

or g=9

OpenStudy (anonymous):

Thanks:)

OpenStudy (anonymous):

2g+14+5g = 8g -3+8 7g + 14 = 8g +5 14=g+5 g= 9 Looks like ive been beat to it :PP

OpenStudy (anonymous):

haha you have competition now my friend.